OpenStudy (anonymous):

"Calculate the following improper integrals. If the integral diverges, show why".. I'm struggling to understand this last part of the question. What does one find if an integral does diverge and how does one go about showing it if so?

OpenStudy (anonymous):

what integral?

OpenStudy (anonymous):

If the limit does not exist or it is infinite, then we say that the improper integral is divergent.

OpenStudy (anonymous):

There's three that are given: \[\int\limits_{1}^{\infty}g^{-4}dg\] \[\int\limits_{1}^{2}\frac{ 1 }{ (2-1)^{2} }ds\] \[\int\limits_{0}^{\infty}\frac{ -2t }{e ^{2t} }dt\]

OpenStudy (anonymous):

take the anti derivative, then compute the limit

OpenStudy (anonymous):

Ah, so the second integral, which I found to equal -1/3 + 1/3(infinity) or just infinity, is divergent?

OpenStudy (anonymous):

the second integral has a typo in it

OpenStudy (anonymous):

apologies, the 2 should be a s, yes

OpenStudy (anonymous):

In the denominator

OpenStudy (anonymous):

did you find the anti - derivative?

OpenStudy (anonymous):

i get \[\frac{1}{1-x}\] at \(2\) you have no problem but at \(1\) you are out of luck

OpenStudy (anonymous):

if you want to be more formal \[\lim_{l\to 1}\frac{1}{1-l}\] does not exist

OpenStudy (anonymous):

Perfect - thank you for your help with that!

OpenStudy (anonymous):

you good with the others?

OpenStudy (anonymous):

\[\int x^{-4}dx=-\frac{1}{3x^3}\] and so \[\int_1^{\infty}x^{-4}dx=\lim_{l\to \infty}-\frac{1}{3l^3}+\frac{1}{3}=0+\frac{1}{3}=\frac{1}{3}\]

OpenStudy (anonymous):

Yes thanks - got that one okay! Then used int. by parts for the third which seemed to work